Problem 2 (20 points) A single-angle tension member (L7x4x5/8) has to ension member (L7x4x5/8) has two gage lines in it
Answer #1

a) Gross aree Lexux5/8 1/2*2%24 Jo 0 0 0 el q Ag = 8 (9-1/2) = + (art/2) (+) 2 B 1272 T Age (7-5716) (%) +4-5/16) (54) Рее 6-